OpenStudy (anonymous):

What is value of Pi f(x)= 0 - L

OpenStudy (anonymous):

\[\frac{\pi }{4}=\left(1-\frac{1}{3}+\frac{1}{5}-\frac{1}{7}+\text{...}\text{..}\right)\]

OpenStudy (pokemon23):

I give up I don't know how to do that?

OpenStudy (anonymous):

\[\pi =4\left(1-\frac{1}{3}+\frac{1}{5}-\frac{1}{7}+\text{...}\text{..}\right)\]

OpenStudy (anonymous):

\[\left(1-\frac{1}{3}+\frac{1}{5}-\frac{1}{7}+\text{...}\text{..}\right)=\underset{n=0}{\overset{\infty }{\sum }} (-1)^n \frac{1}{2n+1}\]

OpenStudy (anonymous):

There you have it, exact value of Pi \[4\sum _{n=0}^{\infty } (-1)^n\frac{1}{2n+1}=\pi\]

OpenStudy (pokemon23):

ok what grade do i learn this

OpenStudy (anonymous):

should have learned it in daycare

OpenStudy (pokemon23):

what! didn't go to day care Darn I missed the freaking lesson

OpenStudy (turingtest):

very interesting, but where did the original series come from?

OpenStudy (anonymous):

That's fourier series representation of this piecewise function f(x)= 0 - L<x<0 L 0<x<L

OpenStudy (turingtest):

oh, is that all? I need to review Fourier series. *bookmark

OpenStudy (anonymous):

@bahrom7893

OpenStudy (bahrom7893):

dude i suck at series, and calc 2 and 3 in general.

OpenStudy (bahrom7893):

MATH PROCESSING ERROR

OpenStudy (anonymous):

just playing